A singly ionized helium atom is in the ground state. It absorbs energy and makes a transition to the n = 5 excited state. The ion returns to the ground state by emitting FOUR photons ONLY. What is the wavelength of the second highest energy photon?
Answer #1

Emn= (13.6Z2)((1/m2)-(1/n2)) eV

For transition of second highest energy from 5th energy level putting n= 5 and m= 2 we get

E25 = (13.6 × 42) ×21/25= 182.78 eV

Wavelength of photon = velocity of light / frequency of photon

= hc/E25 = 6.62×10-34 × 3× 108 /( 182.78 × 1.6 × 10-19)

= 0.06791 × 10-7m

= 679.1 × 10-10m